# Integral representation of solutions to initial-boundary value problems in the framework of the hyperbolic heat equation

### v1

@@ Abstract
We consider initial-boundary value problems on a finite interval for the system of the energy balance equation and modified Fourier’s law (constitutive equation, commonly called the Cattaneo equation) describing temperature (or internal energy) and heat flux. This hyperbolic system of first-order partial differential equations is the simplest model of non-Fourier heat conduction. Boundary conditions comprise various models of behavior of a physical system at the boundaries, including boundary conditions describing Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case the boundary conditions express the relationship of unknown functions (temperature or internal energy and heat flux) with each other. To solve the problems, we apply the Fokas unified transform method.
Keywords: non-Fourier heat conduction, Cattaneo equation, modified Fourier’s law, hyperbolic heat equation, initial-boundary value problems, Fokas unified transform method
@@ 1 Introduction
The heat equation
$\partial_{t}T-\alpha\Delta T=0,$ (1.1)
where $T$ is temperature, and $\alpha$ is thermal diffusivity, is commonly used for description of heat conduction in homogeneous isotropic media. The heat equation is obtained from the energy balance equation
$C_{\mathrm{v}}\partial_{t}T+\nabla\cdot\boldsymbol{q}=0$ (1.2)
and Fourer’s law (the constitutive equation)
$\boldsymbol{q}=-\kappa\nabla T,$ (1.3)
@@
where $C_{\mathrm{v}}$ is volumetric heat capacity, $\boldsymbol{q}$ is the vector of heat flux, and $\kappa$ is thermal conductivity. Hence the thermal diffusivity is given by $\alpha=\kappa/C_{\mathrm{v}}$.
The energy balance equation (1.2) exactly expresses the first law of thermodynamics as applied to heat transfer: the conservation of thermal energy. Fourier’s law, in turn, is approximate relation between the heat flux and temperature. Fourier’s law can be considered as a phenomenological relation. In classical irreversible (weakly non-equilibrium or linear non-equilibrium) thermodynamics Fourier’s law is derived under the assumption of local thermodynamic equilibrium [1, 2, 3]. However, this assumption is violated in very small dimensions and short timescales, and at low temperatures [4, 5, 6, 7, 8, 9, 10, 11]. Besides, the heat equation violates the causality principle due to infinite speed of heat propagation in this model.
The simplest and apparently most “famous” modification of Fourer’s law is given by the constitutive equation
@@
$\tau\partial_{t}\boldsymbol{q}+\boldsymbol{q}=-\kappa\nabla T,$ (1.4)
where $\tau$ is the relaxation time. This equation is usually attributed to Cattaneo [4] or even to Maxwell, Cattaneo and Vernotte [12, 13]. However, Cattaneo’s priority is questionable [5, 14], see also Ref.[15]. Therefore, we will call this equation modified Fourier’s law. The constitutive equation (1.4) is the first and simplest constitutive equation obtained in the framework of various extended thermodynamics: rational extended thermodynamics [16], extended irreversible thermodynamics [17] and non-equilibrium thermodynamics with internal variables [18], see also Ref. [19].
@@
It follows from the energy balance equation (1.2) and the modified Fourier’s law (1.4) that temperature satisfies the so called hyperbolic heat equation
$\tau\partial_{t}^{2}T+\partial_{t}T-\alpha\Delta T=0.$ (1.5)
Since this equation is hyperbolic, it provides a finite speed of heat propagation and therefore satisfies the causality principle.
@@
Solving initial value problems (in the entire space) for this and other equations obtained in the framework of extended thermodynamics and analysing their solutions is straightforward [20, 21, 22]. However, solving initial-boundary value problems (IBVPs) analytically may encounter difficulties. Below, for simplicity, we will consider equations in one-dimensional space. Also, we will consider problems on a finite interval. Problems on the half line can be solved in the same way.
In one-dimensional space the hyperbolic heat equation (1.5) takes the form
$\tau\partial_{t}^{2}T+\partial_{t}T-\alpha\partial_{x}^{2}T=0.$ (1.6)
@@
$\tau\partial_{t}^{2}q+\partial_{t}q-\alpha\partial_{x}^{2}q=0.$ (1.7)
@@
However, other boundary conditions may be set as well, e. g., one or both boundary conditions can describe Newton’s law (convection boundary conditions), which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case IBVPs cannot be formulated for either equation (1.6) or equation (1.7). In this case IBVPs should be formulated for the system of the energy balance equation (1.2) and the constitutive equation (1.4) (both in one dimension if we stay in one-dimensional space). And in this case, the Fourier method is not applicable. Application of the Laplace transform to these problems is problematic as well. All these problems can be solved by the Fokas unified transform method (UTM) [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. In this paper we solve these IBVPs by the UTM. The UTM was applied to systems of equations in Refs. [26, 30, 31, 33, 35, 34]. The peculiarity of the problems considered in this paper is that the boundary conditions generally express the relationship of unknown functions (temperature and heat flux) with each other.
In the papers [37, 38, 13, 39, 40] the IBVPs in the framework of the hyperbolic heat equation were solved both analytically and numerically for particular boundary conditions.
@@ 2 Statement of the initial-boundary value problem
We consider the IBVP for the system of the energy balance equation (including the heat source term) and modified Fourier’s law on the interval $(0,l)$. For convenience, we consider internal energy $e=C_{\mathrm{v}}T$ instead of temperature $T$. Therefore, we solve the system of the equations
@@
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0,$ (2.1b)
where $f$ is the heat source term. The system (2.1) is hyperbolic. Initial conditions for the system (2.1) set the initial state of the physical system and are given by
@@
where $|\gamma_{0}|+|\delta_{0}|>0$ and $|\gamma_{l}|+|\delta_{l}|>0$. The general form of the boundary conditions allows us to describe different behavior of the physical system at the boundary: we can set the temperature or heat flux at the boundary; these boundary conditions also include Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. We assume that the initial and boundary conditions are consistent with each other if necessary.
@@
$\varLambda(k)=\begin{pmatrix}0&\mathrm{i}k\\[2.0pt] \mathrm{i}k\beta&1/\tau\end{pmatrix},\quad\beta=\frac{\alpha}{\tau}.$
@@ 3.1 The local relations
First we need to find dispersion relations. It is necessary to note here that the dispersion relations used in the unified transform method differ by a factor of “$\mathrm{i}$” (imaginary unit) from the standard use. We substitute the vector $\boldsymbol{u}=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx-\omega(k)t}\boldsymbol{c}$, where $\boldsymbol{c}$ is an arbitrary constant vector, into the homogeneous system (2.4). Since $\boldsymbol{c}$ is an arbitrary vector, $\omega(k)$ must satisfy the equation
@@
$\omega^{2}-\frac{1}{\tau}\omega+\frac{\alpha}{\tau}k^{2}\equiv\omega^{2}-\frac{1}{\tau}\omega+\beta k^{2}=0.$
@@
$\omega_{1,2}(k)=\frac{1\pm\sqrt{1-4\alpha\tau k^{2}}}{2\tau},$ (3.1)
@@
In general, the matrix operator $X$ is differential, but this is not the case in this problem. Besides, it does not depend on $k$, and is given by
$X=\begin{pmatrix}0&-1\\ -\beta&0\end{pmatrix}.$
@@ 3.2 The global relations
$\boldsymbol{a}(k,t)$ $=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)X\boldsymbol{u}(0,s)\mathop{}\!\mathrm{d}s=\begin{pmatrix}-\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{q}^{(1)}(0,t)\\[3.0pt] \mathrm{i}k\beta\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{q}^{(2)}(0,t)\end{pmatrix},$ (3.8a)
$\boldsymbol{b}(k,t)$ $=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)X\boldsymbol{u}(l,s)\mathop{}\!\mathrm{d}s=\begin{pmatrix}-\mathrm{i}k\beta\tilde{e}^{(1)}(l,t)+\omega_{2}\tilde{q}^{(1)}(l,t)\\[3.0pt] \mathrm{i}k\beta\tilde{e}^{(2)}(l,t)-\omega_{1}\tilde{q}^{(2)}(l,t)\end{pmatrix},$ (3.8b)
the functions $\tilde{e}^{(m)}$ and $\tilde{q}^{(m)}$ are time transforms of the boundary values of $e$ and $q$, respectively, defined by [^1] This time transform differs from the time transform, used in the UTM literature. However, this form of the transform seems to be more convenient.
$\tilde{e}^{(m)}(x,t)\equiv\tilde{e}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}e(x,s)\mathop{}\!\mathrm{d}s,$ (3.9a)
$\tilde{q}^{(m)}(x,t)\equiv\tilde{q}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}q(x,s)\mathop{}\!\mathrm{d}s,$ (3.9b)
(note that $\tilde{e}^{(m)}$ and $\tilde{q}^{(m)}$ do not change under the transformation $k\to-k$), $\tilde{\boldsymbol{F}}$ is expressed through the Fourier transform of $\boldsymbol{F}$ by
$\tilde{\boldsymbol{F}}(k,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s,$
@@
$\hat{f}$ is the Fourier transform of $f$. $\tilde{\boldsymbol{F}}$ can be easily expressed through the Fourier transform of $f$:
$\tilde{\boldsymbol{F}}(k,t)\equiv\begin{pmatrix}\tilde{F}_{1}(k,t)\\[2.0pt] \tilde{F}_{2}(k,t)\end{pmatrix}=\int_{0}^{t}\begin{pmatrix}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}(t-s)}\\[2.0pt] \omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}(t-s)}\end{pmatrix}\hat{f}(k,s)\mathop{}\!\mathrm{d}s.$ (3.11)
@@ 3.3 A preliminary expression for the solution
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}(k,t)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}k(l-x)}S(k)\boldsymbol{b}(k,t)\mathop{}\!\mathrm{d}k.$ (3.12)
This formula is not yet the final representation of the solution to the problem (2.1)–(2.3). Indeed, the formula contains the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The vectors depend on the functions $\tilde{e}^{(1,2)}$ and $\tilde{q}^{(1,2)}$ (see Eq. (3.8)) that are time transforms of the boundary values of the unknown functions $e$ and $q$ (see Eq. (3.9)). Depending on the coefficients in the boundary conditions (2.3), some of these boundary functions or all of them are unknown and need to be found.

### v2

# Integral representation of solutions to initial-boundary value problems in the framework of the Guyer–Krumhansl heat equation

@@ Abstract
We consider initial-boundary value problems (IBVPs) on a finite interval for the system of the energy balance equation and Guyer–Krumhansl constitutive equation. Boundary conditions comprise various models of behavior of a physical system at the boundaries, including boundary conditions describing Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case the boundary conditions express the relationship of unknown functions (temperature or internal energy and heat flux) with each other. To solve the problems, we apply the Fokas unified transform method. To illustrate the final formulae, we consider a numerical example for a special case in which heat exchange at one of the boundaries obeys Newton’s law.
@@ 1 Introduction
The Guyer–Krumhansl heat equation is obtained from the energy balance equation and Guyer–Krumhansl constitutive equation. The energy balance equation is expressed by the equation
$C_{\mathrm{v}}\partial_{t}T+\nabla\cdot\boldsymbol{q}=f(\boldsymbol{r},t),$ (1.1)
@@
where $T\equiv T(\boldsymbol{r},t)$ is temperature, $C_{\mathrm{v}}$ is volumetric heat capacity, $\boldsymbol{q}\equiv\boldsymbol{q}(\boldsymbol{r},t)$ is the vector of heat flux, $f$ is the heat source term.
The Guyer–Krumhansl constitutive equation, relating heat flux and temperature, is given by [1, 2, 3]
$\partial_{t}\boldsymbol{q}+\frac{1}{\tau_{\mathrm{R}}}\boldsymbol{q}+\frac{c_{\mathrm{s}}^{2}C_{\mathrm{v}}}{3}\nabla T=\frac{\tau_{\mathrm{N}}c_{\mathrm{s}}^{2}}{5}\left(\Delta\boldsymbol{q}+2\nabla\nabla\cdot\boldsymbol{q}\right),$ (1.2)
where $c_{\mathrm{s}}$ is the Debye (average) speed of phonons, $\tau_{\mathrm{N}}$ is a relaxation time for momentum-conserving collisions (normal (N-)processes) and $\tau_{\mathrm{R}}$ is a relaxation time for momentum-nonconserving collisions (resistive (R-)processes) in the phonon gas. The Guyer–Krumhansl constitutive equation (1.2) can be written as
@@
$\tau\partial_{t}\boldsymbol{q}+\boldsymbol{q}+\kappa\nabla T=\ell^{2}\left(\Delta\boldsymbol{q}+2\nabla\nabla\cdot\boldsymbol{q}\right),$ (1.3)
where
$\tau=\tau_{\mathrm{R}},\quad\kappa=\frac{\tau_{\mathrm{R}}c_{\mathrm{s}}^{2}C_{\mathrm{v}}}{3}\equiv\frac{\ell^{2}C_{\mathrm{v}}}{3\tau_{\mathrm{R}}},\quad\ell^{2}=\frac{\tau_{\mathrm{R}}\tau_{\mathrm{N}}c_{\mathrm{s}}^{2}}{5},$
$\tau$ is the relaxation time, $\kappa$ is the thermal conductivity, $\ell=c_{\mathrm{s}}\tau$ is the mean free path of the heat carriers (phonons).
@@
It follows from the energy balance equation (1.1) and the Guyer–Krumhansl constitutive equation (1.3) that temperature satisfies the equation
$C_{\mathrm{v}}\left[\tau\partial_{t}^{2}T+\partial_{t}T-\mu^{2}\partial_{t}\Delta T-\alpha\Delta T\right]=\tau\partial_{t}f+f-\mu^{2}\Delta f.$ (1.4)
where
$\alpha=\frac{\kappa}{C_{\mathrm{v}}},\quad\mu^{2}=3\ell^{2},$
$\alpha$ is thermal diffusivity.
In the absence of the heat source, i. e., if $f=0$, Eq. (1.4) takes the form
$\tau\partial_{t}^{2}T+\partial_{t}T-\mu^{2}\partial_{t}\Delta T-\alpha\Delta T=0,$ (1.5)
This is the Guyer–Krumhansl heat equation, as it is usually written [2].
The Guyer–Krumhansl model was applied to heat conduction in Refs. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
Solving initial value problems (in the entire space) for this and other equations obtained in the framework of extended thermodynamics and analysing their solutions is straightforward [15, 16, 17]. In the papers [6, 7, 8, 10, 11, 13] initial-boundary value problems (IBVPs) for the Guyer–Krumhansl heat equation were solved both analytically and numerically for particular boundary conditions. However, solving IBVPs analytically in general case may encounter difficulties.
Below, for simplicity, we will consider equations in one-dimensional space.
In one-dimensional space Eq. (1.5) takes the form
$\tau\partial_{t}^{2}T+\partial_{t}T-\mu^{2}\partial_{t}\partial_{x}^{2}T-\alpha\partial_{x}^{2}T=0.$ (1.6)
@@
$\tau\partial_{t}^{2}q+\partial_{t}q-\mu^{2}\partial_{t}\partial_{x}^{2}q-\alpha\partial_{x}^{2}q=0.$ (1.7)
@@
However, other boundary conditions may be set as well, e. g., one or both boundary conditions can describe Newton’s law (convection boundary conditions), which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case IBVPs cannot be formulated for either equation (1.6) or equation (1.7). In this case IBVPs should be formulated for the system of the energy balance equation (1.1) and Guyer–Krumhansl constitutive equation (1.3) (both in one dimension if we stay in one-dimensional space). And in this case, the Fourier method is not applicable. Application of the Laplace transform to these problems is problematic as well. All these problems can be solved by the Fokas unified transform method (UTM) [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. In this paper we solve these IBVPs by the UTM. The UTM was applied to systems of equations in Refs. [21, 25, 26, 28, 30, 29]. The peculiarity of the problems considered in this paper is that the boundary conditions generally express the relationship of unknown functions (temperature and heat flux) with each other.
@@ 2 Statement of the initial-boundary value problem
We consider the IBVP for the system of the energy balance equation (including the heat source term) and Guyer–Krumhansl constitutive equation on the interval $(0,l)$. For convenience, we consider internal energy $e=C_{\mathrm{v}}T$ instead of temperature $T$. Therefore, we solve the system of the equations
@@
$\tau\partial_{t}q+q-\mu^{2}\partial_{x}^{2}q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0,$ (2.1b)
where $f$ is the heat source term. Initial conditions for the system (2.1) set the initial state of the physical system and are given by
@@
where $|\gamma_{0}|+|\delta_{0}|>0$ and $|\gamma_{l}|+|\delta_{l}|>0$.
@@
$\varLambda(k)=\begin{pmatrix}0&\mathrm{i}k\\[5.0pt] \dfrac{\alpha}{\tau}\mathrm{i}k&\dfrac{\mu^{2}}{\tau}k^{2}+\dfrac{1}{\tau}\end{pmatrix}\equiv\begin{pmatrix}0&\mathrm{i}k\\[5.0pt] \mathrm{i}k\beta&\theta k^{2}+\dfrac{1}{\tau}\end{pmatrix}$.
@@ 3.1 The local relations
First we need to find dispersion relations. They are determined by the equation
@@
$\omega^{2}-\frac{1+\mu^{2}k^{2}}{\tau}\omega+\frac{\alpha}{\tau}k^{2}=0.$
@@
$\omega_{1,2}(k)=\frac{1+\mu^{2}k^{2}\pm\sqrt{(1+\mu^{2}k^{2})^{2}-4\alpha\tau k^{2}}}{2\tau},$ (3.1)
@@
In this problem
$X(k)=\begin{pmatrix}0&-1\\[3.0pt] -\dfrac{\alpha}{\tau}&\dfrac{\mu^{2}}{\tau}(\mathrm{i}k+\partial_{x})\end{pmatrix}\equiv\begin{pmatrix}0&-1\\[2.0pt] -\beta&\theta(\mathrm{i}k+\partial_{x})\end{pmatrix},\quad\beta=\dfrac{\alpha}{\tau},\quad\theta=\dfrac{\mu^{2}}{\tau}.$
@@ 3.2 The global relations
$\boldsymbol{a}(k,t)$ $=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)X\boldsymbol{u}(0,s)\mathop{}\!\mathrm{d}s$ (3.8a)
$=\begin{pmatrix}-\big[\mathrm{i}k\beta\tilde{e}_{0}^{(1)}-(\omega_{2}-\theta k^{2})\tilde{q}_{0}^{(1)}-\mathrm{i}k\theta\tilde{q}_{1}^{(1)}\big](0,t)\\[3.0pt] \big[\mathrm{i}k\beta\tilde{e}_{0}^{(2)}-(\omega_{1}-\theta k^{2})\tilde{q}_{0}^{(2)}-\mathrm{i}k\theta\tilde{q}_{1}^{(2)}\big](0,t)\end{pmatrix}$,
$\boldsymbol{b}(k,t)$ $=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)X\boldsymbol{u}(l,s)\mathop{}\!\mathrm{d}s$ (3.8b)
$=\begin{pmatrix}-\big[\mathrm{i}k\beta\tilde{e}_{0}^{(1)}-(\omega_{2}-\theta k^{2})\tilde{q}_{0}^{(1)}-\mathrm{i}k\theta\tilde{q}_{1}^{(1)}\big](l,t)\\[3.0pt] \big[\mathrm{i}k\beta\tilde{e}_{0}^{(2)}-(\omega_{1}-\theta k^{2})\tilde{q}_{0}^{(2)}-\mathrm{i}k\theta\tilde{q}_{1}^{(2)}\big](l,t)\end{pmatrix}$,
the functions $\tilde{e}_{j}^{(1,2)}$ and $\tilde{q}_{j}^{(1,2)}$ are time transforms of $e$ and $q$ and the derivatives $\partial_{x}e$ and $\partial_{x}q$, respectively, defined by
$\tilde{e}_{j}^{(m)}(x,t)\equiv\tilde{e}_{j}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}\partial_{x}^{j}e(x,s)\mathop{}\!\mathrm{d}s,$ (3.9a)
$\tilde{q}_{j}^{(m)}(x,t)\equiv\tilde{q}_{j}(\omega_{m}(k),x,t)=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\omega_{m}(k)(t-s)}\partial_{x}^{j}q(x,s)\mathop{}\!\mathrm{d}s,$ (3.9b)
(note that $\tilde{e}_{j}^{(m)}$ and $\tilde{q}_{j}^{(m)}$ do not change under the transformation $k\to-k$), $\tilde{\boldsymbol{F}}$ is expressed through the Fourier transform of $f$ by
$\tilde{\boldsymbol{F}}(k,t)\equiv\begin{pmatrix}\tilde{F}_{1}(k,t)\\[2.0pt] \tilde{F}_{2}(k,t)\end{pmatrix}=\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varOmega(k)(t-s)}S^{-1}(k)\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\\ =\int_{0}^{t}\begin{pmatrix}-\omega_{2}\mathop{\mathrm{e}}\nolimits^{-\omega_{1}(t-s)}\\[2.0pt] \omega_{1}\mathop{\mathrm{e}}\nolimits^{-\omega_{2}(t-s)}\end{pmatrix}\hat{f}(k,s)\mathop{}\!\mathrm{d}s,$ (3.10)
where
@@
$\hat{f}$ is the Fourier transform of $f$.
@@ 3.3 A preliminary expression for the solution
$\boldsymbol{u}(x,t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)t}\hat{\boldsymbol{u}}_{0}(k)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\int_{0}^{t}\mathop{\mathrm{e}}\nolimits^{-\varLambda(k)(t-s)}\hat{\boldsymbol{F}}(k,s)\mathop{}\!\mathrm{d}s\mathop{}\!\mathrm{d}k\\ -\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}(k,t)\mathop{}\!\mathrm{d}k+\frac{1}{2\pi}\int_{-\infty}^{\infty}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}(k,t)\mathop{}\!\mathrm{d}k.$ (3.12)
This formula is not yet the final representation of the solution to the problem (2.1)–(2.3). Indeed, the formula contains the vectors $\boldsymbol{a}$ and $\boldsymbol{b}$. The vectors depend on the functions $\tilde{e}_{0}^{(1,2)}$ and $\tilde{q}_{0,1}^{(1,2)}$ (see Eq. (3.8)) that are time transforms of the boundary values of the unknown functions $e$ and $q$ and their spatial derivatives (see Eq. (3.9)). Depending on the coefficients in the boundary conditions (2.3), some of these boundary functions or all of them are unknown and need to be found.